OpenStudy (anonymous):

Factor completely 49x2 – 81

OpenStudy (cj7529):

49x^2 - 81 49 = 7^2 81 = 9^2 so, 49x^2 - 81 =(7x + 9)(7x - 9)

OpenStudy (anonymous):

Thanks!